OpenStudy (anonymous):

Write the equation so that y is a function of x. 1.x+7y=0 2.3x=2y-18 3.4y-x=20-y

OpenStudy (angela210793):

1)y=-x/7 2) y=3x/2+9 3) y=x/5+4

OpenStudy (anonymous):

any reason they don't say "solve for y"?

OpenStudy (anonymous):

sattelite can you solve the one before this